Operation Summary for August 10, 2011 - August 12, 2011 |
Wednesday August 10 The day shift began with Operations monitoring the Tevatron (TEV) store 8966 (75.54E30) and the Recycler stash (158.0E10), with the Antiproton Source (Pbar) stacking (58.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBFT-1010 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Linac personnel investigated a ground water leak into the 750 KeV area. A Linac quadrupole power supply (L:QPS115) tripped its breaker for the first of three times in nine minutes; operators and Linac technicians investigated. Operators searched and secured MI-31 at 9:58 AM. MI personnel began coalescing studies. Operations terminated store 8966 at 11:34 AM and put the TEV into shot setup. Operators prepared the Meson M01 enclosure for access at 11:45 AM to look at the alarming sump pump (M01SP5) and, during the Meson downtime, EE Support personnel changed the PLC for a Switchyard power supply (HP3DS). Operations established store 8967 at 12:53 PM with and average initial luminosity (AIL) of 127.63E30. Linac personnel began a couple of study periods, beginning at 1:03 PM, which interrupted beam for a very short amount of time. During the evening shift, Operations disabled beam to T-1010 at 6:07 PM. Operators searched and secured MI-31; Pelletron personnel began the Electron Cooling checkout, turn on, and conditioning at 8:21 PM. Operations tried to terminate store 8967 at 10:50 PM, but there was a problem moving collimators to safe positions; operators contacted TEV personnel and, with their help, temporarily resolved the problem. Operations terminated store 8967 at 11:34 PM, but continued to investigate the collimator problem. Thursday August 11 The midnight shift began with TEV collimator investigations continuing, with Operations monitoring the stash (141.1E10), with Pbar stacking (186.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators reported that the collimator issue was resolved at 12:49 AM. Operators put the TEV into shot setup at 12:59 AM and established store 8968 at 2:59 AM with an AIL of 132.58E30. At 7:22 AM, a MI RF station (MIRF10) tripped off and would not reset; RF technicians investigated and then took it offline. Pelletron conditioning continued all shift. During the day shift, Operations terminated store 8968 at 8:30 AM and put the TEV into shot setup. At 8:46 AM, MIRF11 tripped off repeatedly; a RF tech investigated. Operations established store 8969 at 9:48 AM with an AIL of 152.73E30. A RF tech reported that MIRF10 & 11 had failing modulators and it would take about 35 minutes for repairs; Operations turned off NuMI for the work. Operators reported at 11:10 AM that Pbar develop problems while trying to receive reverse protons from MI; Pbar personnel investigated, adjusted the scale factors in the P1 transfer line and adjusted a current setting for a power supply (HT701). Operations established beam to T-1010 at 12:40 PM. MI personnel began a coalescing study at 1:46 PM. During the evening shift, Operations terminated store 8969 at 5:43 PM. At 5:57 PM, operators reported that the TEV had locked up a 90 GeV; operators investigated and found a bad uninterruptable power supply (UPS) at TEV sector A2. An EE Support person came in and had the UPS changed out by 7:05 PM. Operations put the TEV through a dry squeeze at 7:48 PM followed by a wet squeeze at 8:29 PM. Operations disabled beam to T-1010 at 9:09 PM. Operations put the TEV into shot setup at 9:16 PM. Operators reported that the Pelletron kept tripping off and recovering; operators talked with a Pelletron person who suggested an adjustment to the corona event aggregate, which solved the problem. Operations established store 8973 at 10:29 PM with an AIL of 235.6E30. Operators reported short downtimes due to problems with a Linac Klystron RF station (KRF1), a Linac quadrupole power supply (L:QPS108), BRF3, BRF7, BRF8, BRF9, and BRF17. Friday August 12 The midnight shift began with Operations monitoring store 8973 (161.83E30) and the stash (124.5E10), with Pbar stacking (20.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 1:03 AM, a TEV B2 service building air conditioner tripped off for the first of many times in one hour and 40 minutes; a FESS person investigated and added some Freon to the system. At 4:13 AM, BRF12 tripped off and would not reset; operators bypassed the station. Operators tuned the 400 MeV line at 6:02 AM. Operators conducted a Booster injection tune at 6:23 AM. Operations established beam to T-1010 at 7:13 AM. At 7:15 AM, RF personnel reported a LCW leak in MIRF6 station; they investigated, reported a cracked hose barb, and had it repaired 30 minutes later. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 08/12/2011 email Fermilab |
| Security, Privacy, Legal | |